# Comparison of Extended Lubrication Theories for Stokes Flow

### v1

@@ 5 Conclusions
We considered several models in extended lubrication theory for an incompressible, low Reynolds number fluid across a range of geometries, including several in which the standard assumptions of classical lubrication theory are not strictly satisfied. The VA-ELT model proposed here amends the velocity and pressure of the T.G.-ELT solution to ensure that incompressibility and the boundary conditions are satisfied. The solutions of velocity and pressure for the Reynolds, VA-ELT, T.G.-ELT, $\varepsilon^{k}$-PLT and Stokes models were compared for the logistic step and the triangular slider.
We found that all lubrication models have increased error in pressure and velocity as the magnitude of surface variation increases, while the length scale ration $\varepsilon=L_{x}/L_{y}$ is kept constant. In the presence of a large surface gradient, the Stokes solution exhibits significant $\frac{\partial p}{\partial y}$ in a region surrounding the surface gradient, compared with the ELT and $\varepsilon^{k}$-PLT solutions which localize this pressure variation to regions where the surface gradient is large. The consequences of this localization is overestimation of $\frac{\partial p}{\partial y}$, overestimation of the velocity magnitude, and in the case of very large surface gradients, spurious flow recirculation.
For the logistic step, the VA-ELT and $\varepsilon^{k}$-PLT solutions are an improvement on the Reynolds solution for smaller slopes; but at larger slopes, particularly for $\varepsilon^{k}$-PLT, the error exceeds that of Reynolds equation. For smaller slopes, the VA-ELT solution has the smallest error in velocity, and the $\varepsilon^{2}$-PLT solution has the smallest error in pressure. At larger slopes, the VA-ELT and $\varepsilon^{k}$-PLT solutions exhibit flow recirculation at too small of slope compared with the Stokes solution, and the velocity magnitude in the recirculation region becomes unreasonably large given the small Reynolds number.
@@ Appendix A The Reynolds equation finite difference solution
Define the uniform discretisation of $x_{0},x_{L}]$,
(31) $\{x_{i}\}_{i=0}^{N},$ $x_{i}=x_{0}+i\Delta x,$ $\Delta x=\tfrac{1}{N}|x_{L}-x_{0}|.$
@@
(32) $\frac{1}{2\Delta x^{2}}\bigg(\Big(h_{i+1}^{3}+h_{i}^{3}\Big)p_{i+1}-\Big(h_{i+1}^{3}+2h_{i}^{3}+h_{i-1}^{3}\Big)p_{i}+\Big(h_{i}^{3}+h_{i-1}^{3}\Big)p_{i-1}\bigg)\\ =\frac{6\eta\mathcal{U}}{2\Delta x}\Big(h_{i+1}-h_{i-1}\Big).$
@@ Appendix B The Stokes Equations
The biharmonic formulation of the incompressible Navier-Stokes equations is an effective method of solution for low Reynolds number flows [7, 9, 12, 3]. Through introduction of the stream function $\psi(x,y)$ satisfying,
@@
When $Re=0$, the velocity-stream formulation Eq. 34 reduces to the biharmonic Stokes equation $\nabla^{4}\psi=0$. The characteristic pressure used to formulate Eq. 34 is $P_{*}^{\prime}=\varepsilon P_{*}=\eta U_{*}/L_{y}$, and thus the Stokes model is not sensitive to surface variation as the lubrication models are.
To compare the Stokes solution with the one dimensional lubrication theory solution, we assume the inlet and outlet flow profiles correspond to a fully developed laminar flow with flux $\mathcal{Q}$. The surface boundary conditions for velocity are Eqs. 8 and 9, and the pressure satisfies Eq. 10. The inlet and outlet velocity profiles are,
(35) $u(x_{0},y)$ $=u_{\text{Re}}(x_{0},y)$ $\frac{\partial u}{\partial x}\Big|_{x_{L},y}=0,$
(36) $v(x_{0},y)$ $=0$ $v(x_{L},y)=0,$
where $u_{\text{Re}}(x,y)$ is the Reynolds equation velocity Eq. 11 expressed in terms of $\mathcal{Q}$ from Eq. 13. The corresponding boundary conditions for the stream function $\psi$ are,
(37) $\psi(x_{0},y)$ $=\int_{0}^{y}u_{\text{Re}}(x,\hat{y})d\hat{y}$ $\frac{\partial\psi}{\partial x}\Big|_{x_{L},y}=0,$
(38) $\psi(x,0)$ $=0$ $\psi(x,h(x))=\mathcal{Q},$
@@
(39) $\int_{0}^{y}u_{\text{Re}}(x,\hat{y})d\hat{y}=\frac{\mathcal{Q}y^{2}}{\big[h(x)\big]^{3}}\Big(3h(x)-2y\Big)+\frac{\mathcal{U}y}{\big[h(x)\big]^{2}}\Big(h(x)-y\Big)^{2}.$
The solution $\psi$, $u$, $v$ to the biharmonic Navier-Stokes equations Eqs. 34 and 33 is determined through an iterative second-order accurate finite difference method as given in [3, 7].
@@
(40) $\frac{\partial p}{\partial x}\Big|_{i,j}=\frac{\nu}{\Delta x^{2}}\Big(u_{i-1,j}+u_{i+1,j}-4u_{i,j}+u_{i,j-1}+u_{i,j+1}\Big)\\ -\frac{1}{2\Delta x}\Big(u_{i,j}(u_{i+1,j}-u_{i-1,j})+v_{i,j}(u_{i,j+1}-u_{i,j-1})\Big),$
and $\frac{\partial p}{\partial y}$ as in Eq. 2 is similar. The pressure $p(x,y)$ is then determined numerically by a path integral from the outlet.
@@ Appendix C Perturbed Lubrication Theory
(41) $P_{2}(\bar{x},\bar{y})=-\frac{\partial U_{0}}{\partial\bar{x}}+\int_{0}^{\bar{x}}\frac{d\gamma_{3}}{d\hat{\bar{x}}}d\hat{\bar{x}}$,
(42) $U_{2}(\bar{x},\bar{y})=\frac{d^{2}\big[\bar{h}^{-3}\big]}{d\bar{x}^{2}}\Big(\bar{y}^{4}-\bar{h}^{3}\bar{y}\Big)-2\frac{d^{2}\big[\bar{h}^{-2}\big]}{d\bar{x}^{2}}\Big(\bar{y}^{3}-\bar{h}^{2}\bar{y}\Big)+\frac{1}{2}\frac{d\gamma_{3}}{d\bar{x}}(\bar{y}^{2}-\bar{h}\bar{y}),$
@@
(43) $V_{2}(\bar{x},\bar{y})=-\frac{d^{3}\big[\bar{h}^{-3}\big]}{d\bar{x}^{3}}\Big(\tfrac{1}{5}\bar{y}^{5}-\tfrac{1}{2}\bar{h}^{3}\bar{y}^{2}\Big)+\frac{3}{2}\frac{d^{2}\big[\bar{h}^{-3}\big]}{d\bar{x}^{2}}\frac{d\bar{h}}{d\bar{x}}\bar{h}^{2}\bar{y}^{2}\\ +\frac{d^{3}\big[\bar{h}^{-2}\big]}{d\bar{x}^{3}}\Big(\tfrac{1}{2}\bar{y}^{4}-\bar{h}^{2}\bar{y}^{2}\Big)-2\frac{d^{2}\big[\bar{h}^{-2}\big]}{d\bar{x}^{2}}\frac{d\bar{h}}{d\bar{x}}\bar{h}\bar{y}^{2}\\ -\frac{1}{2}\frac{d^{2}\gamma_{3}}{d\bar{x}^{2}}\Big(\tfrac{1}{3}\bar{y}^{3}-\tfrac{1}{2}\bar{h}\bar{y}^{2}\Big)+\frac{1}{4}\frac{d\gamma_{3}}{d\bar{x}}\frac{d\bar{h}}{d\bar{x}}\bar{y}^{2},$
@@
(44) $\frac{d\gamma_{3}}{d\bar{x}}=-\frac{18}{5}\frac{d^{2}\big[\bar{h}^{-3}\big]}{d\bar{x}^{2}}\bar{h}^{2}+6\frac{d^{2}\big[\bar{h}^{-2}\big]}{d\bar{x}^{2}}\bar{h}.$
@@
(45) $P_{4}(\bar{x},\bar{y})=-\frac{\partial U_{2}}{\partial\bar{x}}+\int_{0}^{\bar{y}}\frac{\partial^{2}V_{0}}{\partial{\bar{x}^{2}}}d\hat{\bar{y}}+\int_{0}^{\bar{x}}\frac{d\gamma_{5}}{d\hat{\bar{x}}}d\hat{\bar{x}}$,
(46) $U_{4}(\bar{x},\bar{y})=-\frac{1}{20}\frac{d^{4}\big[\bar{h}^{-3}\big]}{d\bar{x}^{4}}\Big(\bar{y}^{6}-\bar{h}^{5}\bar{y}\Big)+\frac{3}{20}\frac{d^{4}\big[\bar{h}^{-2}]}{d\bar{x}^{4}}\Big(\bar{y}^{5}-\bar{h}^{4}\bar{y}\Big)\\ +\Bigg(\frac{1}{3}\frac{d^{2}\phi_{1}}{d\bar{x}^{2}}-\frac{2}{3}\frac{d^{2}\phi_{2}}{d\bar{x}^{2}}+\frac{1}{6}\frac{d^{2}\phi_{3}}{d\bar{x}^{2}}\Bigg)\Big(\bar{y}^{3}-\bar{h}^{2}\bar{y}\Big)\\ -\frac{1}{12}\frac{d^{3}\gamma_{3}}{d\bar{x}^{3}}\Big(\bar{y}^{4}-\bar{h}^{3}\bar{y}\Big)+\frac{1}{2}\frac{d\gamma_{5}}{d\bar{x}}\Big(\bar{y}^{2}-\bar{h}\bar{y}\Big),$
@@
(47) $V_{4}(\bar{x},\bar{y})=-\frac{1}{20}\frac{d^{5}\big[\bar{h}^{-3}\big]}{d\bar{x}^{5}}\Big(\tfrac{1}{7}\bar{y}^{7}-\tfrac{1}{2}\bar{h}^{5}\bar{y}^{2}\Big)+\frac{1}{8}\frac{d^{4}\big[\bar{h}^{-3}\big]}{d\bar{x}^{4}}\frac{d\bar{h}}{d\bar{x}}\bar{h}^{4}\bar{y}^{2}\\ +\frac{3}{20}\frac{d^{5}\big[\bar{h}^{-2}\big]}{d\bar{x}^{5}}\Big(\tfrac{1}{6}\bar{y}^{6}-\tfrac{1}{2}\bar{h}^{4}\bar{y}^{2}\Big)-\frac{3}{10}\frac{d^{4}\big[\bar{h}^{-2}\big]}{d\bar{x}^{4}}\frac{d\bar{h}}{d\bar{x}}\bar{h}^{3}\bar{y}^{2}\\ +\frac{1}{3}\Bigg(\frac{d^{3}\phi_{1}}{d\bar{x}^{3}}-2\frac{d^{3}\phi_{2}}{d\bar{x}^{3}}\Bigg)\Big(\tfrac{1}{4}\bar{y}^{4}-\tfrac{1}{2}\bar{h}^{2}\bar{y}^{2}\Big)-\Bigg(\frac{d^{2}\phi_{1}}{d\bar{x}^{2}}-2\frac{d^{2}\phi_{2}}{d\bar{x}^{2}}\Bigg)\frac{d\bar{h}}{d\bar{x}}\bar{h}\bar{y}^{2}\\ +\frac{1}{6}\Bigg(\frac{d^{3}\phi_{3}}{d\bar{x}^{3}}\Big(\tfrac{1}{4}\bar{y}^{4}-\tfrac{1}{2}\bar{h}^{2}\bar{y}^{2}\Big)-\frac{d^{2}\phi_{3}}{d\bar{x}^{2}}\frac{d\bar{h}}{d\bar{x}}\bar{h}\bar{y}^{2}\Bigg)-\frac{1}{12}\frac{d^{4}\gamma_{3}}{d\bar{x}^{4}}\Big(\tfrac{1}{5}\bar{y}^{5}-\tfrac{1}{2}\bar{h}^{3}\bar{y}^{2}\Big)\\ +\frac{1}{8}\frac{d^{3}\gamma_{3}}{d\bar{x}^{3}}\frac{d\bar{h}}{d\bar{x}}\bar{h}^{2}\bar{y}^{2}+\frac{1}{2}\frac{d^{2}\gamma_{5}}{d\bar{x}^{2}}\Big(\tfrac{1}{3}\bar{y}^{3}-\tfrac{1}{2}\bar{h}\bar{y}^{2}\Big)-\frac{1}{4}\frac{d\gamma_{5}}{d\bar{x}}\frac{d\bar{h}}{d\bar{x}}\bar{y}^{2},$
@@
(48) $\frac{d^{2}\phi_{1}}{d\bar{x}^{2}}=\frac{d^{4}\big[\bar{h}^{-3}]}{d\bar{x}^{4}}\bar{h}^{3}+6\frac{d^{3}\big[\bar{h}^{-3}\big]}{d\bar{x}^{3}}\frac{d\bar{h}}{d\bar{x}}\bar{h}^{2}+\frac{d^{2}\big[\bar{h}^{-3}\big]}{d\bar{x}^{2}}\Bigg(6\bigg[\frac{d\bar{h}}{d\bar{x}}\bigg]^{2}\bar{h}+3\frac{d^{2}\bar{h}}{d\bar{x}^{2}}\bar{h}^{2}\Bigg),$
(49) $\frac{d^{2}\phi_{2}}{d\bar{x}^{2}}=\frac{d^{4}\big[h^{-2}\big]}{d\bar{x}^{4}}\bar{h}^{2}+4\frac{d^{3}\big[\bar{h}^{-2}\big]}{d\bar{x}^{3}}\frac{d\bar{h}}{d\bar{x}}\bar{h}+2\frac{d^{2}\big[\bar{h}^{-2}\big]}{d\bar{x}^{2}}\Bigg(\bigg[\frac{d\bar{h}}{d\bar{x}}\bigg]^{2}+\frac{d^{2}\bar{h}}{d\bar{x}^{2}}\bar{h}\Bigg),$
(50) $\frac{d^{2}\phi_{3}}{d\bar{x}^{2}}=\frac{d^{3}\gamma_{3}}{d\bar{x}^{3}}\bar{h}+2\frac{d^{2}\gamma_{3}}{d\bar{x}^{2}}\frac{d\bar{h}}{d\bar{x}}+\frac{d^{2}\bar{h}}{d\bar{x}^{2}}\frac{d\gamma_{3}}{d\bar{x}},$
@@
(51) $\frac{d\gamma_{5}}{d\bar{x}}=\frac{3}{14}\frac{d^{4}\big[\bar{h}^{-3}\big]}{d\bar{x}^{4}}\bar{h}^{4}-\frac{3}{5}\frac{d^{4}\big[\bar{h}^{-2}]}{d\bar{x}^{4}}\bar{h}^{3}+\frac{3}{10}\frac{d\gamma_{3}}{dx}\bar{h}^{2}.$

### v2

@@ 5 Conclusions
We considered several models in extended lubrication theory and compared the solutions of velocity and pressure across a range of geometries, including several configurations in which the thin film assumption of classical lubrication theory is not strictly satisfied. The VA-ELT solution proposed here amends the velocity and pressure of the T.G.-ELT solution to ensure that incompressibility and the boundary conditions are satisfied. To assess the sensitivity of the lubrication models to the thin film assumption and to large magnitudes of surface variation in the low Reynolds number limit, the Reynolds, VA-ELT, T.G.-ELT, $\varepsilon^{2}$-PLT and $\varepsilon^{4}$-PLT solutions were compared with the Stokes solution for the logistic step, the backward facing step, and the triangular slider with positive and negative texturing.
We found that all lubrication models have increased error in pressure and velocity as the magnitude of surface variation increases, while the length scale ratio $\varepsilon=L_{y}/L_{x}$ is kept constant. In the presence of a large surface gradient, the Stokes solution exhibits significant $\frac{\partial p}{\partial y}$ in a region surrounding the surface gradient, compared with the ELT and $\varepsilon^{k}$-PLT solutions which localize this pressure variation to regions where the surface gradient is large. The consequences of this localization is overestimation of $\frac{\partial p}{\partial y}$, overestimation of the velocity magnitude, and in the case of very large surface gradients, spurious flow recirculation.
For the logistic step, the VA-ELT and $\varepsilon^{k}$-PLT solutions are an improvement on the Reynolds solution for smaller slopes; but at larger slopes, particularly for $\varepsilon^{k}$-PLT, the error exceeds that of the Reynolds solution. For smaller slopes, the VA-ELT solution has the smallest error in velocity, and the $\varepsilon^{2}$-PLT solution has the smallest error in pressure. At larger slopes, the VA-ELT and $\varepsilon^{k}$-PLT solutions exhibit flow recirculation at too small of slope compared with the Stokes solution, and the velocity magnitude in the recirculation region becomes unreasonably large given the small Reynolds number.
@@ Appendix A The Reynolds equation finite difference solution
Define the uniform discretisation of $0,1]=([x_{0},x_{L}]-x_{0})/L_{x}$,
(32) $\{X_{i}\}_{i=0}^{N},$ $X_{i}=i\Delta X,$ $\Delta X=\tfrac{1}{N}.$
@@
(33) $\Big(H_{i+1}^{3}+H_{i}^{3}\Big)P_{i+1}-\Big(H_{i+1}^{3}+2H_{i}^{3}+H_{i-1}^{3}\Big)P_{i}+\Big(H_{i}^{3}+H_{i-1}^{3}\Big)P_{i-1}\\ =6U_{b}\Big(H_{i+1}-H_{i-1}\Big)\Delta X.$
@@ Appendix B The Stokes Equations
The biharmonic formulation of the incompressible Navier-Stokes equations is an effective method of solution for low Reynolds number flows [gupta_new_2005, marner_potential-velocity_2014, sen_4oec_2015, biswas_hoc_2017]. Through introduction of the stream function $\Psi(X,Y)$ satisfying,
@@
When $\text{Re}=0$, the velocity-stream formulation Eq. 35 reduces to the biharmonic Stokes equation $\nabla^{4}\Psi=0$. Note, the characteristic pressure used to formulate Eq. 35 is $\tilde{P_{*}}=\varepsilon^{2}P_{*}=\eta U_{*}/L_{x}$, and thus the Stokes model is not sensitive to surface variation as the lubrication models are.
To compare the Stokes solution with the one dimensional lubrication theory solution, we assume the inlet and outlet flow profiles correspond to a fully developed laminar flow with flux $\mathcal{Q}$. The surface boundary conditions for velocity are Eqs. 6 and 7, and the pressure satisfies Eq. 15. The inlet and outlet velocity profiles are,
(36) $U(0,Y)$ $=U_{0}(0,Y)$ $\frac{\partial U}{\partial X}\bigg|_{1,Y}=0,$
(37) $V(0,Y)$ $=0$ $V(1,Y)=0.$
The corresponding boundary conditions for the stream function $\Psi$ are,
(38) $\Psi(0,Y)$ $=\int_{0}^{Y}U_{0}(X,\hat{Y})d\hat{Y}$ $\frac{\partial\Psi}{\partial X}\Big|_{1,Y}=0,$
(39) $\Psi(X,0)$ $=0$ $\Psi(X,H)=1,$
@@
(40) $\int_{0}^{Y}U_{0}(X,\hat{Y})d\hat{Y}=\frac{Y^{2}}{H^{3}}\Big(3H-2Y\Big)+\frac{U_{b}Y}{H^{2}}\Big(H-Y\Big)^{2}.$
The solution $\Psi$, $U$, $V$ to the biharmonic Navier-Stokes equations Eqs. 35 and 34 is determined through an iterative second-order accurate finite difference method as given in [biswas_hoc_2017, gupta_new_2005].
@@
(41) $\frac{\partial P}{\partial X}\Big|_{i,j}=\frac{1}{\Delta X^{2}}\Big(U_{i-1,j}+U_{i+1,j}-4U_{i,j}+U_{i,j-1}+U_{i,j+1}\Big)\\ -\frac{1}{2\Delta X}\Big(U_{i,j}(U_{i+1,j}-U_{i-1,j})+V_{i,j}(U_{i,j+1}-U_{i,j-1})\Big),$
and $\frac{\partial P}{\partial Y}$ as in Eq. 2 is similar. The pressure $P(X,Y)$ is then determined numerically by a path integral from the outlet with the boundary condition $P(1,0)=0$. The average pressure drop $\Delta P_{\text{Stokes}}$ is determined by numerical integration according to Eq. 9 for a given $H$.
@@ Appendix C Perturbed Lubrication Theory
(42) $P_{2}(X,Y)=-\frac{\partial U_{0}}{\partial X}+\int_{0}^{X}\frac{d\gamma_{3}}{d\hat{X}}d\hat{X}$,
(43) $U_{2}(X,Y)=\frac{d^{2}\big[H^{-3}\big]}{dX^{2}}\Big(Y^{4}-H^{3}Y\Big)-2\frac{d^{2}\big[H^{-2}\big]}{dX^{2}}\Big(Y^{3}-H^{2}Y\Big)+\frac{1}{2}\frac{d\gamma_{3}}{dX}(Y^{2}-HY),$
@@
(44) $V_{2}(X,Y)=-\frac{d^{3}\big[H^{-3}\big]}{dX^{3}}\Big(\tfrac{1}{5}Y^{5}-\tfrac{1}{2}H^{3}Y^{2}\Big)+\frac{3}{2}\frac{d^{2}\big[H^{-3}\big]}{dX^{2}}\frac{dH}{dX}H^{2}Y^{2}\\ +\frac{d^{3}\big[H^{-2}\big]}{dX^{3}}\Big(\tfrac{1}{2}Y^{4}-H^{2}Y^{2}\Big)-2\frac{d^{2}\big[H^{-2}\big]}{dX^{2}}\frac{dH}{dX}HY^{2}\\ -\frac{1}{2}\frac{d^{2}\gamma_{3}}{dX^{2}}\Big(\tfrac{1}{3}Y^{3}-\tfrac{1}{2}HY^{2}\Big)+\frac{1}{4}\frac{d\gamma_{3}}{dX}\frac{dH}{dX}Y^{2},$
@@
(45) $\frac{d\gamma_{3}}{dX}=-\frac{18}{5}\frac{d^{2}\big[H^{-3}\big]}{dX^{2}}H^{2}+6\frac{d^{2}\big[H^{-2}\big]}{dX^{2}}H.$
@@
(46) $P_{4}(X,Y)=-\frac{\partial U_{2}}{\partial X}+\int_{0}^{Y}\frac{\partial^{2}V_{0}}{\partial{X^{2}}}d\hat{Y}+\int_{0}^{X}\frac{d\gamma_{5}}{d\hat{X}}d\hat{X}$,
(47) $U_{4}(X,Y)=-\frac{1}{20}\frac{d^{4}\big[H^{-3}\big]}{dX^{4}}\Big(Y^{6}-H^{5}Y\Big)+\frac{3}{20}\frac{d^{4}\big[H^{-2}]}{dX^{4}}\Big(Y^{5}-H^{4}Y\Big)\\ +\Bigg(\frac{1}{3}\frac{d^{2}\phi_{1}}{dX^{2}}-\frac{2}{3}\frac{d^{2}\phi_{2}}{dX^{2}}+\frac{1}{6}\frac{d^{2}\phi_{3}}{dX^{2}}\Bigg)\Big(Y^{3}-H^{2}Y\Big)\\ -\frac{1}{12}\frac{d^{3}\gamma_{3}}{dX^{3}}\Big(Y^{4}-H^{3}Y\Big)+\frac{1}{2}\frac{d\gamma_{5}}{dX}\Big(Y^{2}-HY\Big),$
@@
(48) $V_{4}(X,Y)=-\frac{1}{20}\frac{d^{5}\big[H^{-3}\big]}{dX^{5}}\Big(\tfrac{1}{7}Y^{7}-\tfrac{1}{2}H^{5}Y^{2}\Big)+\frac{1}{8}\frac{d^{4}\big[H^{-3}\big]}{dX^{4}}\frac{dH}{dX}H^{4}Y^{2}\\ +\frac{3}{20}\frac{d^{5}\big[H^{-2}\big]}{dX^{5}}\Big(\tfrac{1}{6}Y^{6}-\tfrac{1}{2}H^{4}Y^{2}\Big)-\frac{3}{10}\frac{d^{4}\big[H^{-2}\big]}{dX^{4}}\frac{dH}{dX}H^{3}Y^{2}\\ +\frac{1}{3}\Bigg(\frac{d^{3}\phi_{1}}{dX^{3}}-2\frac{d^{3}\phi_{2}}{dX^{3}}\Bigg)\Big(\tfrac{1}{4}Y^{4}-\tfrac{1}{2}H^{2}Y^{2}\Big)-\Bigg(\frac{d^{2}\phi_{1}}{dX^{2}}-2\frac{d^{2}\phi_{2}}{dX^{2}}\Bigg)\frac{dH}{dX}HY^{2}\\ +\frac{1}{6}\Bigg(\frac{d^{3}\phi_{3}}{dX^{3}}\Big(\tfrac{1}{4}Y^{4}-\tfrac{1}{2}H^{2}Y^{2}\Big)-\frac{d^{2}\phi_{3}}{dX^{2}}\frac{dH}{dX}HY^{2}\Bigg)-\frac{1}{12}\frac{d^{4}\gamma_{3}}{dX^{4}}\Big(\tfrac{1}{5}Y^{5}-\tfrac{1}{2}H^{3}Y^{2}\Big)\\ +\frac{1}{8}\frac{d^{3}\gamma_{3}}{dX^{3}}\frac{dH}{dX}H^{2}Y^{2}+\frac{1}{2}\frac{d^{2}\gamma_{5}}{dX^{2}}\Big(\tfrac{1}{3}Y^{3}-\tfrac{1}{2}HY^{2}\Big)-\frac{1}{4}\frac{d\gamma_{5}}{dX}\frac{dH}{dX}Y^{2},$
@@
(49) $\frac{d^{2}\phi_{1}}{dX^{2}}=\frac{d^{4}\big[H^{-3}]}{dX^{4}}H^{3}+6\frac{d^{3}\big[H^{-3}\big]}{dX^{3}}\frac{dH}{dX}H^{2}+\frac{d^{2}\big[H^{-3}\big]}{dX^{2}}\Bigg(6\bigg[\frac{dH}{dX}\bigg]^{2}H+3\frac{d^{2}H}{dX^{2}}H^{2}\Bigg),$
(50) $\frac{d^{2}\phi_{2}}{dX^{2}}=\frac{d^{4}\big[h^{-2}\big]}{dX^{4}}H^{2}+4\frac{d^{3}\big[H^{-2}\big]}{dX^{3}}\frac{dH}{dX}H+2\frac{d^{2}\big[H^{-2}\big]}{dX^{2}}\Bigg(\bigg[\frac{dH}{dX}\bigg]^{2}+\frac{d^{2}H}{dX^{2}}H\Bigg),$
(51) $\frac{d^{2}\phi_{3}}{dX^{2}}=\frac{d^{3}\gamma_{3}}{dX^{3}}H+2\frac{d^{2}\gamma_{3}}{dX^{2}}\frac{dH}{dX}+\frac{d^{2}H}{dX^{2}}\frac{d\gamma_{3}}{dX},$
@@
(52) $\frac{d\gamma_{5}}{dX}=\frac{3}{14}\frac{d^{4}\big[H^{-3}\big]}{dX^{4}}H^{4}-\frac{3}{5}\frac{d^{4}\big[H^{-2}]}{dX^{4}}H^{3}+\frac{3}{10}\frac{d\gamma_{3}}{dx}H^{2}.$